Altman Corporation has interest expenses of $120,000 annually. Altman’s annual sales are $4 million, its tax rate is 25%, and its net profit margin on sales is 10 percent. What is Altman’s TIE?

1
Expert's answer
2021-11-22T13:24:59-0500

TIE Ratio = Earnings Before Interest and Taxes/Interest Expenses.

EBIT = Operating Expenses - Gross Profit

=120000 - (10/100*4000000)

= 120000 - 40000

= $80000

TIE Ratio = 120000/80000

= 1.5X

Therefore; TIE = 1.5X